OpenStudy (anonymous):

The sides of triangle have measures of x,x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences, in order based on the question. Be sure your equations are fully simplified.

OpenStudy (maheshmeghwal9):

y is 4cm longer than x so y will be equal to x only when u add 4cm to x So \[\color{red}{y=x+4.}\]

OpenStudy (maheshmeghwal9):

& all the sides of the triangle sum up-to 46cm. So; \[\color{red}{x+x+y=46}\]\[\color{blue}{\implies x+x+(x+4)=46.}\]It from first eq. So u gt\[3x=42.\]

OpenStudy (maheshmeghwal9):

so \[\huge{x= ?}\]

OpenStudy (anonymous):

x=14

OpenStudy (maheshmeghwal9):

\[\color{blue}{??????????????????????????????????????????????????????????????????????????????????????????????????????}\]